# Iterative assembly of ${}^{171}$b atom arrays in cavity-enhanced optical lattices

### v2

@@ III Repeated loading to build up a large array
In figure 3b, the atom number is inferred from the image taken before rearrangement, so the measured fill fraction is sensitive to all losses that occur between such images. This fill fraction is typically 98%, with the largest single loss contribution coming from vacuum loss during the typical 300 ms time between images due to our approximately 30 s vacuum lifetime. We can add additional diagnostic images directly after rearrangement to determine the atom number available for computation, and observe typical filling fractions of 99%, again dominated by vacuum loss during the 150 ms typically allocated for rearrangement. Our experimental control allocates a variable duration for rearrangement that depends on the number and length of moves to be performed, as well as the time required to compute the moves. The durations above are typical for the 1225-site arrays studied in this work. Other sources of loss or potential loss are described in more detail in Appendix B, and include the handoff between the tweezer arrays and the lattice, and imaging loss and infidelity.
@@ IV Conclusions and outlooks
We have demonstrated an iterative method for assembling large arrays of individually controlled neutral atoms, suitable for quantum computation, simulation, and metrology. Through this approach, we decouple the final size of the array from the number of atoms that can be loaded at once, and the need to create many tweezer traps with sufficient depth for loading and imaging atoms. In particular, our reservoir and lattice traps are much deeper than the target tweezer traps, allowing for the use of large numbers of sites with moderate laser power. Our current arrays have percent-level defects, which could be straightforwardly reduced with improved vacuum.
For quantum computing applications, our iterative loading may be combined with recently demonstrated mid-circuit measurement techniques [25, 26] in order to achieve continuous mid-circuit refilling of array defects. This capability will be critical for mitigating the effect of atom loss during execution of complex error-corrected circuits that may extend beyond the lifetime of an individual atom in the array.
V Appendix A: Raman Sideband Cooling
@@
Figure 4: Raman sideband cooling (RSC). (a) Diagram of relevant directions and polarizations. RSC beams RSC1 and RSC2 are combined to cool the $z$ direction, while RSC1 and RSC3 are combined to cool the $x$ and $y$ directions. Optical pumping (beam OP) is applied along the $x$ direction, which is along the applied magnetic field. (b) Level diagram for Raman sideband cooling. Raman transitions detuned by 35 MHz from ${}^{3}$${}_{1}$ $F=3/2$ $m_{f}=1/2$ are driven from ${}^{1}$${}_{0}$ $m_{f}=1/2$ to ${}^{1}$${}_{0}$ $m_{f}=-1/2$, while reducing the number of motional quanta along the addressed direction(s) from $n$ to $n-1$. Optical pumping through ${}^{3}$${}_{1}$ $F=1/2$ $m_{f}=1/2$ returns atoms to ${}^{1}$${}_{0}$ $m_{f}=1/2$ to enable further cycles. (c) Sideband spectra in the $x$, $y$ plane. The sideband imbalance indicates an average number of motional quanta along the differential momentum vector of the RSC1 and RSC3 beams of $\bar{n}_{xy}=0.08(1)$. (d) Sideband spectra in the $z$ direction. The sideband imbalance indicates an average number of motional quanta along the $z$ direction of $\bar{n}_{z}=0.12(1)$. (e) Heating in the lattice, as measured from the sideband imbalance in the $x$, $y$ plane, with a fit to an exponential growth profile, returning a time-constant of 190(40) ms.
@@
Transferring atoms from the lattice into shallow tweezer traps with low loss requires that the atoms be much colder than the depth of the tweezer traps. By operating with colder atoms, we can use shallower tweezers, and in turn scale to larger array sizes. To accomplish this, we use Raman sideband cooling (RSC) between the two ground nuclear spin states [21, 25]. For the motional Raman transitions, we use two pairs of beams oriented along the $x$ and $y$ directions, and along the $x$ and $z$ directions to provide cooling along all three directions (fig. 4a). The Raman transitions transfer atoms from ${}^{1}$${}_{0}$ $m_{f}=1/2$ to ${}^{1}$${}_{0}$ $m_{f}=-1/2$, state, and are detuned from the motional carrier transition by the appropriate trap oscillation frequency in order to reduce the motional state by one quanta (fig. 4b). Optical pumping is provided by a beam oriented along $x$ and the magnetic field that addresses the ${}^{1}$${}_{0}$ $m_{f}=-1/2$ to ${}^{1}$${}_{1}$ $F=1/2$, $m_{f}=1/2$ transition.
@@
We perform 20 iterations of cooling, with each iteration consisting of a $\pi$ pulse on the red motional sideband for each pair, followed by optical pumping. The pulse durations are 200 $\mu$s for the $xy$ pair and 100 $\mu$s for the $xz$ pair (tuned to the $z$ direction sidband), and 50 $\mu$s for optical pumping. The total cooling sequence lasts 8 ms. In typical sequences, we apply a 2 ms Doppler cooling pulse prior to the RSC sequence using the same beam as for the imaging pulse, but with lower intensity and farther red-detuning. The Doppler cooling results in 10 $\mu$K temperature in the cavity as measured by a release and recapture protocol [37]. This roughly corresponds to 1 motional quanta occupancy in the cavity. Subsequent RSC reduces the average motional quanta to in each direction to $\sim$ 0.1, as evident from the sideband imbalance following cooling (fig. 4c, d).
@@
VI Appendix B: Sources of Per-Cycle Loss

### v3

@@ II Repeated loading to build up a large array
During loading, we only perform imaging prior to rearrangement. These images are required to determine which sites need to be filled, and any additional images would lead to increased per-cycle loss. The atom numbers and fill fractions inferred from these images are represented as black points in figure 2. Because these images are performed prior to rearrangement, the measured fill fraction is sensitive to any losses that have occurred since the equivalent image was performed in the previous loading cycle. This fill fraction is typically 98%, with the largest single loss contribution coming from vacuum loss during the typical 300 ms time between images due to our approximately 30 s vacuum lifetime. Small variations in fill fraction may be attributed to drifting alignment of the tweezer arrays relative to lattices (see appendix H). We can add additional diagnostic images directly after rearrangement to determine the atom number available for computation (Shown as red markers in figure 2b), and observe typical filling fractions of 99%. For these diagnostic images, the imperfect filling is dominated by vacuum loss during the shorter 150 ms interval typically allocated for rearrangement. Other sources of loss or potential loss are described in more detail in appendix F, and include the handoff between the tweezer arrays and the lattice, and imaging loss and infidelity.
The number of atoms that we can currently load is limited by available laser power for the target array traps. For larger tweezer arrays than those used here, the shallower tweezers lead to additional per-cycle loss when transferring atoms from the lattice to tweezers. This reduces the peak fill fraction, and if the array size is increased enough, leads to a situation where the number of atoms transferred from the reservoir on each cycle is insufficient to replace lost atoms. The number of usable traps could be increased through the use of lasers with higher power or wavelengths with higher polarizability, as well as through more efficient optical paths and reduced optical aberrations.
As with all measurement-based rearrangement protocols, our filling fraction is limited by the atom loss that occurs between the image that identifies vacancies and the completion of the subsequent rearrangement step. In our case, this loss is dominated by background gas collisions that occur during rearrangement. This could be improved either through lower vacuum pressure or by shortening the rearrangement sequence. Several options exist for the latter: A second iteration of rearrangement could be applied, condensing atoms into a slightly smaller sub-array [32]. Because the number of vacancies to be filled would be lower and the distance that atoms must be transported to fill the vacancies would be shorter, this would enable a shorter rearrangement period, lower vacuum losses, and higher fill fraction. Additionally, the rearrangement speed itself could be increased by improving optical aberrations on the rearrangement tweezer.
III Trapping and imaging atoms in a 3D cavity enhanced lattice
Figure 3: Cavity-enhanced optical lattices for rapid low-loss imaging. (a) The intersection of two cavity modes provide three-dimensional confinement for atoms within the field of view of our high-numerical-aperture imaging system. Confinement in the $x$ and $y$ directions is provided by a self-intersecting standing-wave cavity, while confinement in the $z$ direction is provided by a separate self-intersecting travelling-wave cavity mode. Imaging light (not shown for clarity) is incident along the mode of the travelling-wave lattice. (b) Observed counts during a 7 ms long image of single atoms within the optical lattice, showing well-resolved peaks for occupied and unoccupied sites. For these conditions, per-image survival is 99.8(1)%. (c) Lattice homogeneity across the target array as characterized by light shifts on the 1S0 $m_{F}=1/2$ to 3P1, $F=3/2$ $m_{F}=-1/2$ transition. Atoms within 13.4 $\mu$m squares are averaged together for statistics and display clarity. (d, e) Light shift contributions from each cavity averaged over rows (d) and columns (e). Red and orange markers and lines represent the measured and predicted profiles of the XY (red) and Z (orange) cavities, with the prediction centered on the array.
Rapid low-loss imaging of the locations of atoms is essential to our repeated loading protocol. Because atoms are heated by the photon scattering required for imaging, optical traps that are much deeper than the equilibrium atomic temperature established during imaging are required to prevent atom loss. Optical tweezers represent a common method for achieving deep optical traps – the tight foci can provide high intensities for moderate powers. However, the total power requirement for a large array of tweezers presents scaling limitations due to limited available laser power. Cavity-enhanced optical lattices offer an alternative path to achieving many deep traps. While the laser intensity in a lattice is spread over a larger cross-sectional area than in optical tweezers both because it spans the regions between utilized sites, and because the Gaussian mode profile must be large compared to the size of the array to achieve uniform trapping, interference of many reflected paths in a cavity-enhanced lattice can enable more power-efficient generation of traps. For our system, the power required to generate the 1225 imaging sites used in this work is roughly two orders of magnitude lower than the power that would be required to create the same number and depth of traps using optical tweezers of the same wavelength (see appendix I). This provides a promising path to scaling to much larger atomic ensembles.
We employ two intersecting optical cavities to provide tight confinement in three dimensions, as shown in figure 3a. The first cavity – which we call the XY cavity – is a retro-reflected four-mirror cavity whose mode intersects itself at the location of the atoms. The polarization is perpendicular to the plane of propagation, providing interference between the two crossed portions of the cavity mode [33]. Compared to using two crossed non-interfering lattices [34, 18], and assuming equal cavity finesse, this configuration provides an eight-fold enhancement in trap depth (defined as the potential barrier between adjacent lattice sites) per unit of total laser power. The topology of this lattice configuration is also stable to translation of any mirror [33], so the enhanced depth comes without penalty with respect to stability, and also benefits from fewer input-coupling paths and laser stabilization systems. The XY cavity provides tight confinement in the directions perpendicular to the high numerical aperture optical axis used for the tweezers and imaging.
The second cavity, which we call the Z cavity, is a four-mirror travelling-wave cavity whose mode intersects itself at the location of the atoms [35] at 15 degree angles from the XY plane. This provides confinement along the tweezer axis. The polarization of the Z cavity is also oriented perpendicular to its intersection plane, enabling complete interference of the crossing modes (like the XY cavity, the contrast of this interference pattern is also robust to the longitudinal position of the mirrors).
Each cavity has two highly reflecting mirrors, and two matched partially transmitting mirrors. The finesse of the XY and Z cavities are 2900(100) and 3000(100) respectively, as extracted from the measured cavity linewidths of 307(10) kHz and 448(15) kHz (photon lifetimes of 518(15) ns and 355(10) ns) and measured cavity free-spectral-ranges of 890.38 MHz and 1345.13 MHz. Our observed finesse is consistent with the specified transmission of our mirror coatings and measurements in a test setup bound per-mirror losses to 50 parts-per-million, indicating a transmission-dominated cavity. The finesses of the cavities were chosen to attain a large power-buildup while limiting conversion of laser frequency noise to amplitude noise that can occur in cavities with narrow linewidth. The design mode waists of the XY, Z cavities at the location of the atoms are 268 $\mu$m and 183 $\mu$m respectively. For imaging, we typically operate with XY (Z) trap frequencies near 160 kHz (50 kHz), corresponding to 330 $\mu$K (260 $\mu$K) deep traps.
We image atoms in in the cavity using the narrow-linewidth (180 kHz) 1S0 to 3P1 transition, which simultaneously provides cooling and scattering of photons for detection [36, 37, 27]. We use a single linearly polarized imaging beam, whose polarization is perpendicular to the magnetic field, and the propagation direction has projection onto both the X and Z directions. The imaging beam is detuned several hundred kHz from the $F=3/2$, $m_{F}=3/2$ state, and has intensity approximately equal to the resonant saturation intensity of the imaging transition. For applications like array assembly where we wish to determine the presence of an atom but not the state of its nuclear spin (the qubit we use for quantum information applications), we simultaneously apply optical pumping on the 1S0 $m_{F}=-1/2$ to 3P1 $F=1/2$, $m_{F}=1/2$ transition with light incident along the magnetic field ($x$ direction in figure 3). At our benchmark imaging parameters, we collect 50 photons in 7 ms, and distinguish occupied from unoccupied sites with a fidelity of $99.92^{+0.03}_{-0.04}$%, determined from the consistency of classification in repeated images (see appendix G). Per-image atom loss in such repeated images is $2(1)\times 10^{-3}$. Spin-selective imaging requires us to omit the optical pumping, and rely on the combination of frequency and polarization selectivity within the 3P1 manifold to create a large imbalance of scattering between the two nuclear spin states, as demonstrated in optical tweezers in refs. [37, 27]. The omission of the optical pumping does not modify the photon scattering rate appreciably, but does introduce a finite spin flip probability of of $4(1)\times 10^{-3}$ for our typical imaging parameters, measured as an additional apparent loss in repeated images of 1S0 $m_{F}=1/2$. These spin flips are likely due to the fact that the lattice polarization is perpendicular to the magnetic field, and so can induce spin-changing transitions within the 3P1 manifold [37].
In order to ensure uniform detuning of the imaging transition across the array and between different motional states, we operate with a trapping wavelength that has equal polarizability for the ground and excited states of the imaging transition. For the 1S0 $m_{F}=1/2$ to 3P1 $m_{F}=3/2$ transition, we find such a condition near a wavelength of 783.8 nm for lattice polarization perpendicular to the quantization axis. For more details, see appendix D.
We characterize the homogeneity of the optical lattice potential using spectroscopy of the 1S0 $m_{F}=1/2$ to 3P1 $F=3/2$, $m_{F}=-1/2$ transition, for which the excited state has approximately 40% higher polarizability than the ground state. We infer the trap depth from the transition light shifts, measured using optical pumping from the 1S0 $m_{F}=1/2$ state, followed by spin-selective imaging (fig. 3c). By varying the power in each cavity, we can extract their independent contributions, displayed averaged over rows and columns in figure  3d,e. Over our 115 $\mu$m square array, peak deviations are within 20% for the Z lattice and 10% for the XY lattice. Because we operate with a magic wavelength condition for critical operations, these deviations have only a small impact on imaging performance – the minimum trap depth defines the limit on imaging speed for the whole array and excess trap depth leads to additional Raman scattering (discussed in appendix F).
@@ IV Conclusions and outlooks
We have demonstrated an iterative method for assembling arrays of over 1200 individually controlled neutral atoms, suitable for quantum computation, simulation, and metrology. Through this approach, we decouple the final size of the atomic ensemble from the number of atoms that can be loaded at once, and the need to create many tweezer traps with sufficient depth for loading and imaging atoms. In particular, our reservoir and lattice traps are much deeper than the target tweezer traps, allowing for the use of large numbers of sites with moderate laser power.
For quantum computing applications, our iterative loading may be combined with recently demonstrated mid-circuit measurement techniques [26, 27] in order to achieve continuous mid-circuit refilling of array defects. In such a protocol, a subset of atoms (“ancilla qubits”) would be measured and potentially replaced while the rest of the atoms (“data qubits”) maintain coherence as required to continue a quantum computation. The roles of data and ancilla qubits may be alternated in order to allow replacement of any atom within the array. In order to protect data qubits from decoherence from scattering imaging photons, tweezers operating at a wavelength that creates large light shifts to the imaging transition (as is the case for our target-array tweezers) can be applied to the locations of the data qubits, while all atoms are held in the lattice and illuminated with imaging light. This shifts the data qubits off resonance with the imaging light, preventing decoherence, as we have demonstrated in a similar system with tweezer-confined atoms [27]. Mid-circuit reloading also requires that light scattered from the MOT does not cause decoherence of the data qubits. We have demonstrated that in our system, operating a broad-line MOT causes only a 0.03(2)/s decoherence rate for atoms in the science region [27]. We expect that the core-shell configuration used here will have an even lower decoherence rate, due to a substantially reduced scattering rate (appendix A). By combining these capabilities with the continuous loading approach demonstrated here, it should be possible to enable the replacement of lost atoms within a circuit, which is a key capability for the execution of complex error-corrected circuits that may extend beyond the lifetime of an individual atom in the array. For such a protocol to be successful, it is beneficial for all operations performed on atoms to have low rates of atom loss. This is typically required for high-fidelity coherent operations, and with the techniques presented in this work and in reference [27], is possible for atomic readout as well.
Other approaches for mid-circuit measurement with neutral atoms include the use of two atomic species [29], a readout zone [12], an optical cavity [38], or shelving of population to a state that does not couple to imaging light with either alkali [39] or alkaline-earth [26] atoms. Extending these mid-circuit measurement protocols to mid-circuit reloading will face additional challenges. If the reloading is to be performed quasi-indefinitely, there must be a way to replenish a depleted reservoir while maintaining coherence in the remaining atoms. This has been demonstrated using the dual species approach [28], though not with other methods. As we describe above, our approach of separating the MOT loading region from the science region should also be compatible with mid-circuit reservoir refilling and we have previously demonstrated mid-circuit reloading from a reservoir in [27].
Unlike the dual species method, where the process of reloading ancilla atoms into the tweezers leads to the loss of atoms of that species already present, our approach requires only a small fraction of atoms to be reloaded on each reloading cycle. This may ultimately provide an important advantage for mid-circuit reloading in systems with large numbers of atoms.
Appendices
Figure 4: (a) Timing diagram for our continuous loading sequence. See main text and appendices A-E for a detailed description of the referenced operations. The beam paths referenced in the call-out are illustrated in figure 5. The following acronyms are used in the figure: MOT: magneto-optical trap, LAC: light-assisted collisions, OP: optical pumping, RSC: Raman sideband cooling.
Appendices A-H describe Specific elements of our loading sequence. An overview of the sequence with timing information can be found in figure 4. Appendix I provides a quantitative comparison of the power requirements for creating a given array of traps with optical tweezers versus cavity-enhanced lattices.
Appendix A Core-shell MOT
In order to minimize the potential for light scattered from the MOT to cause heating or decoherence of atoms in the science region, we operate with a core-shell configuration [31]. In this MOT, each beam consists of a ring of 399 nm light near resonance with the broad linewidth 1S0 to 1P1 transition. This surrounds a central region with only 556 nm light that is near resonance with the narrow linewidth 1S0 to 3P1 transition. To create collimated hollow beams while minimizing the loss of laser power (compared to blocking the center), we use diffractive (HOLO/OR, DA-069-399-Y-A)-refractive (Thorlabs, AX252-A) axicon pairs. The diffractive axicon eliminates the round-tip effect seen with refractive axicons [40]. The MOT is loaded from a precooled and collimated atomic beam provided by a commercial source (AOSense, Inc.). The shell is primarily responsible for capturing atoms from the incident beam, while the low Doppler temperature of the narrow transition used in the core enables cooling to several 10 $\mu$K. Because atoms do not scatter light from the broad transition except during initial capture, the core-shell design poses reduced risk of generating scattered photons that may be absorbed by atoms in the target array – the total scattering rate on the narrow transition is much lower than the broad transition, and the large magnetic field present in the science region dramatically suppresses reabsorption of photons via the narrow-line transition. In our operating condition, the 399 nm photon scattering rate is about two orders of magnitude lower than the broad-line stage of a sequential MOT. We find that the core-shell MOT configuration increases the MOT loading rate by roughly a factor of two (to $1.5\times 10^{6}$ atoms per second), compared to a more standard sequential MOT configuration in a static magnetic field. At this loading rate, we can collect enough atoms to saturate the filling of our reservoir within 80 ms. This allows us to overlap the MOT loading in time with other operations performed in the science region.
Appendix B Transport Lattice and loading of reservoir array
From the core-shell MOT, atoms are loaded into a vertically oriented standing-wave lattice formed by counterpropagating 532 nm laser beams focused at the location of the MOT with a beam diameter (1/e2) of 75 $\mu$m. The typical depth of the lattice is 30 $\mu$K. Loading is achieved by overlapping the lattice with the center of the MOT, followed by reducing the detuning and intensity of the 556 nm light. We then extinguish the MOT light and transport the atoms vertically by synchronously applying a frequency offset to the upward-going and downward-going lattice beams, and translating their foci over 100 ms with a profile calculated to minimize jerk [27]. This brings the atoms 30 cm vertically into the vicinity of the target array, with several tens of thousands of atoms transported in a typical cycle. The typical survival during transport is 70 %.
We transfer the atoms from the transport lattice into a reservoir array formed by 483 nm tweezers (a magic wavelength for the 1S0 to 3P1 transition [27]). To avoid disturbing atoms already loaded into the target array, this transfer takes place at a location displaced 170 $\mu$m horizontally from the target array. No dissipation is required to transfer atoms from the transport lattice to the tweezers – loading is achieved by simply increasing the depth of the tweezers from zero to 250 $\mu$K over 1 ms while the lattice is overlapped and then decreasing the intensity of the transport lattice to zero over 1 ms. Once loaded, the reservoir array is translated over 10 ms to be directly adjacent to the target array by changing the angle of a galvo mirror that is imaged onto the entrance pupil of the objective. Compared to acousto-optic deflectors, the galvo mirror provides higher efficiency and a larger scan angle.
Appendix C Target tweezer arrays
The target tweezer array consists of a rectangular grid of tweezer spots formed with light at 459.5960(5) nm wavelength, a magic wavelength for the 1S0 to 3P0 clock transition [27, 41], and a set of tweezer spots formed by light at 423.31 nm, which is near resonance with a transition between 3P1 to a higher-lying 6s8s 3S1 state. All tweezers have waists of approximately 500 nm. Both wavelengths provide large light-shifts to atoms in 3P1, which prevents unwanted scattering from atoms loaded in these tweezer arrays, and can be used interchangeably for this work. For this work, we operate with a 35x35-site target array with 3.3 $\mu$m spacing, comprised of two side-by-side rectangular arrays each using one of the two previously mentioned wavelengths. The typical trap depth in the target array is approximately 50 $\mu$K, which shifts 3P1 by approximately 9 MHz in the 459 nm traps, and 50 MHz in the 423 nm traps (though this can be changed dramatically by varying the detuning from the 3P1 to 3S1 transition). This requires approximately 200 $\mu$W delivered to the atoms per tweezer for the 459 nm traps and 100 $\mu$W per tweezer for the 423 nm traps. We have also operated with overlapped full-size arrays to achieve sufficient trap depth, and obtained similar results. However, the overlapped condition is more sensitive to alignments, and so we use spatially separated arrays for all results presented here.
While translating the reservoir array adjacent to the target array (described in appendix B), both arrays are illuminated with light resonant with the 1S0 to 3P1 $F=3/2$, $m_{F}=3/2$ transition. This induces light-assisted collisions (LACs) between atoms in the reservoir tweezers, resulting in either zero or one atom in each reservoir tweezer with approximately equal probability [19]. Importantly, the tweezer light illuminating the atoms in the target array prevents them from scattering the light used to induce collisions, which might otherwise cause loss of existing atoms in the target array.
Appendix D Rearrangement from reservoir to target tweezer arrays
The rearrangement process relies on first identifying which sites of the reservoir and target arrays are occupied. This is accomplished by transferring the atoms from the tweezer arrays to the cavity-enhanced lattice, and then imaging the atom locations. Transfer is performed by turning on the lattice and simultaneously turning off the tweezers with linear intensity ramps lasting 1 ms. The imaging leaves the atoms at a temperature of 20 $\mu$K, which is too warm to efficiently transfer back into the relatively shallow tweezers of the target array. Following each image, we apply Doppler and Raman sideband cooling (RSC) in the optical lattice to reach an average of $\sim$ 0.1 motional quantum in each direction (appendix E). The low temperatures obtained by RSC allow the atoms to be transferred efficiently into tweezer arrays of lower depth, which in turn allows for the generation of roughly 10 percent larger numbers of tweezers at fixed laser power compared to Doppler cooling alone. Once cooled, the atoms are transferred back into the tweezer arrays, and a single optical tweezer derived from the same laser as the reservoir light is used to transfer atoms from filled reservoir sites into empty target sites. This rearrangement tweezer is approximately 150 $\mu$K deep, which is chosen to balance the efficiency and speed with which an atom can be moved with disturbance of atoms already loaded into the array (see appendix F and figure 6). Our experimental control allocates a variable duration for rearrangement that depends on the number and length of moves to be performed, as well as the time required to compute the moves. The typical duration for the 1225-site arrays studied in this work is approximately 150 ms. A possible future upgrade could involve using multiple rearrangement tweezers to move atoms in parallel, and thus speed up initial filling of the target array.
Efficient imaging benefits from state-insensitive trapping for the 1S0 to 3P1 $F=3/2$, $m_{F}=3/2$ transition, which we find near 783.8 nm for trap polarization perpendicular to the quantization axis. This polarization-dependent magic condition (the polarizability of the excited state is 56% greater for the orthogonal polarization of light) has a fractional deviation from state-insensitive conditions of 0.007/nm. Because the buildup cavities feature non-normal mirror reflections, they are highly birefringent and so provide polarization filtering, mitigating possible effects from drifting lattice polarization on the imaging transition frequency.
Appendix E Raman Sideband Cooling
@@
Figure 5: Raman sideband cooling (RSC). (a) Diagram of relevant directions and polarizations. RSC beams RSC1 and RSC2 are combined to cool the $z$ direction, while RSC1 and RSC3 are combined to cool the $x$ and $y$ directions. Optical pumping (beam OP) is applied along the $x$ direction, which is along the applied magnetic field. (b) Level diagram for Raman sideband cooling. Raman transitions detuned by 35 MHz from 3P1 $F=3/2$ $m_{F}=1/2$ are driven from 1S0 $m_{F}=1/2$ to 1S0 $m_{F}=-1/2$, while reducing the number of motional quanta along the addressed direction(s) from $n$ to $n-1$. Optical pumping through 3P1 $F=1/2$ $m_{F}=1/2$ returns atoms to 1S0 $m_{F}=1/2$ to enable further cycles. (c) Sideband spectra in the $x$, $y$ plane. The sideband imbalance indicates an average number of motional quanta along the differential momentum vector of the RSC1 and RSC3 beams of $\bar{n}_{xy}=0.08(1)$. (d) Sideband spectra in the $z$ direction. The sideband imbalance indicates an average number of motional quanta along the $z$ direction of $\bar{n}_{z}=0.12(1)$. (e) Heating in the lattice, as measured from the sideband imbalance in the $x$, $y$ plane, with a fit to an exponential growth profile, returning a time-constant of 190(40) ms.
@@
Transferring atoms from the lattice into shallow tweezer traps with low loss requires that the atoms be much colder than the depth of the tweezer traps. To accomplish this, we apply Raman sideband cooling (RSC) between the two ground nuclear spin states [22, 26] before transferring atom from the lattice to the tweezers. For the motional Raman transitions, we use two pairs of beams oriented along the $x$ and $y$ directions, and along the $x$ and $z$ directions to provide cooling along all three directions (fig. 5a). The Raman transitions transfer atoms from 1S0 $m_{F}=1/2$ to 1S0 $m_{F}=-1/2$, state, and are detuned from the motional carrier transition by the appropriate trap oscillation frequency in order to reduce the motional state by one quantum (fig. 5b). Optical pumping is provided by a beam oriented along $x$ and the magnetic field that addresses the 1S0 $m_{F}=-1/2$ to 1P1 $F=1/2$, $m_{F}=1/2$ transition.
@@
We perform 20 iterations of cooling, with each iteration consisting of a $\pi$ pulse on the red motional sideband for each pair, followed by optical pumping. The pulse durations are 200 $\mu$s for the $xy$ pair and 100 $\mu$s for the $xz$ pair (tuned to the $z$ direction sideband), and 50 $\mu$s for optical pumping. The total cooling sequence lasts 8 ms. In typical sequences, we apply a 2 ms Doppler cooling pulse prior to the RSC sequence using the same beam as for the imaging pulse, but with lower intensity and farther red-detuning. The Doppler cooling results in 10 $\mu$K temperature in the cavity as measured by a release and recapture protocol [42]. This roughly corresponds to 1 motional quantum occupancy in the cavity. Subsequent RSC reduces the average motional quanta in each direction to $\sim$ 0.1, as evident from the sideband imbalance following cooling (fig. 5c, d).
@@
Appendix F Sources of Per-Cycle Loss